OpenStudy (anonymous):

Using the square root method... 5x^2+35=0

OpenStudy (amistre64):

you mean completing the square?

OpenStudy (amistre64):

or its kin

OpenStudy (anonymous):

it says solve by the square root method

OpenStudy (amistre64):

5(x^2+7) = 0, there is no real number that will do this

OpenStudy (anonymous):

I know, it will end up with i in the answer

OpenStudy (amistre64):

x^2 + 7 = 0 x^2 = -7 x = +- i sqrt(7)

OpenStudy (anonymous):

Thanks

OpenStudy (amistre64):

i got no idea if that the "square root" method or not tho

OpenStudy (anonymous):

Nope, I know you have to make it 5x^2=-35 and then I dont really know how to proceed. Thanks

OpenStudy (amistre64):

divide off the 5

OpenStudy (amistre64):

and you get to x^2 = -7

OpenStudy (anonymous):

When do you take the sq rt from each side?

OpenStudy (amistre64):

when you get x^2 all alone in order to operate on it

OpenStudy (amistre64):

the sqrt and the ^2 cancel each other out

OpenStudy (amistre64):

if it (5x)^2 , then you can do it as is; but the way you wrote it, you only have an x^2